OpenStudy (anonymous):

If f(x) = 3x – 5, then what is f–1(x) ? (the -1 is suppose to be an exponent)

OpenStudy (anonymous):

Is it supposed to be an inverse equation?

OpenStudy (anonymous):

it means the inverse of f to find inverse you interchange x and y and then make y the subject of the formula

OpenStudy (anonymous):

If so it's (x+5)/3=y

OpenStudy (anonymous):

the choices are x-5/3....x+3/5....x-3/5....x+5/3

OpenStudy (anonymous):

f(x) = 3x-5 y=3x-5 x=3y-5 x+5=3y x/3 +5/3 =y

OpenStudy (anonymous):

which is x+5/3